The most recent financial statements for Assouad, Inc., are shown here: Income Statement Sales Costs Balance Sheet Current $8,700 assets $3,600 Current $ 2,400 liabilities Long-term 6,100 Fixed assets 9,000 3,980 Taxable income $2,600 Equity 6,220 Taxes (24%) 624 Total $12,600 Total $12,600 Net income $1,976 Assets, costs, and current liabilities are proportional to sales. Long-term debt and equity are not. The company maintains a constant 35 percent dividend payout ratio. As with every other firm in its industry, next years sales are projected to increase by exactly 18 percent What is the external financing needed? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) External financing needed
Answer #1
Sales(8700*1.18) 10266
Costs(6100*1.18) 7198
Taxable income $3068
Taxes@24%(3068*0.24) $736.32
Net income $2331.68
Less:dividend($2331.68*35%) $816.088
Addition to retained earnings $1515.592

Total assets would be=$12600*1.18=$14868

Total equity=$6220+Addition to retained earnings

=$6220+1515.592=$7735.592

Current liabilities=(2400*1.18)=$2832

Total assets=Total equity+Total liabilities

Hence external financing needed=$14868-$7735.592-$(2832+3980)

=$320.41(Approx).